(3 points) (a) Find the general solution to y + 2y = 0. Give your answer as y=... . In your answer, use c1 and c2 to denote(3 points) (a) Find the general solution to y′′+2y′=0. Give your answer as y=... . In your answer, use c1 and c2 to denote arbitrary constants and x the independent variable. Enter c1 as c1 and c2 as c2.
Answer #1

(a) To Solve &- y+ay=o - (1) 2 (n) Put n Y=l, y=rqy and you arh then corresponding characteristic equation becomes for UUsing initial value y (0)=1 -200) =) Cit Ca e = 1 a) 6+ C2 = 1 - Q2 (as e 21) By using y(0) = 1 =) 6.-26g 22 (0) - 1 -) C1-2Rules for finding solution for busilary .. - Equation Cor complementary solution case-ls Roots are real and Dietinet 1 Y = clast page represent rule for general solution according to different roots of characteristic equation.